# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 5.2 The correspondence between difference operators and polygons
Suppose $v_{i}\in M$ are such that the sequence $P_{i}=Rv_{i}$ is a polygon. Then there exists a properly bounded degree $d$ difference operator $\mathcal{D}{{}_{v}}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\mathcal{D}{{}_{v}}v=0$.
@@
The following says that the correspondence $v\mapsto\mathcal{D}_{v}$ from Proposition 5.13 provides a right inverse for the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Thus, this map is surjective, and the proof of Theorem 5.7 is complete.
@@
Assume that $v_{i}\in M$, where $M$ is a rank $d$ free left $R$-module, is such that the sequence $Rv_{i}$ is a polygon. Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ be a properly bounded degree $d$ difference operator such that $\mathcal{D}v=0$. Then there is a projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ mapping the polygon associated with $\mathcal{D}$ to the polygon $(Rv_{i})$.
@@
Since $\mathcal{D}v=0$, and the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$, by Proposition 5.6 we have $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. This means that the image of the mapping $\psi\colon M^{*}\to\mathbb{R}^{\mathbb{Z}}$ given by $\xi\mapsto(\xi(v_{i}))$ is precisely $\mathrm{Ker}\,\mathcal{D}$. Furthermore, since $M^{*}$ and $\mathrm{Ker}\,\mathcal{D}$ are free modules of the same rank, and $R$ is stably finite, it follows that $\psi\colon M^{*}\to\mathrm{Ker}\,\mathcal{D}$ is a right $R$-module isomorphism. Consider the dual $\psi^{*}\colon(\mathrm{Ker}\,\mathcal{D})^{*}\to M$. It is a left $R$-module isomorphism, and, for any $\xi\in M^{*}$, we have
@@ 6.1 Pentagram maps in terms of difference operators
By Theorem 5.7, polygons in the (left) projective plane $\mathbb{P}^{2}(R)$, considered up to projective transformation, are in one-to-one correspondence with elements of the set $R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$ of properly bounded left difference operators of degree three, up to left-right action of $(R^{\mathbb{Z}})^{*}$. Accordingly, one can describe the pentagram map as a dynamical system on difference operators. In this section we give an intrinsic description of that dynamical system. For technical reasons, we consider the inverse pentagram map, rather than the pentagram map itself. It turns out that that this inverse, phrased in terms of difference operators, is described by a simple linear equation. In the case when the ground ring is a field, this result was established in [8].
In order to have well-defined dynamics on difference operators, we only consider polygons whose image under the inverse pentagram map is a polygon (so that the image corresponds to a difference operator). Denote the inverse pentagram map by $\Psi$, and let $\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$ be the set of projective equivalence classes of polygons in $\mathbb{P}^{2}(R)$ whose image under the inverse pentagram map is a polygon. That way, we have a map
$\Psi\colon\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}\to\mathfrak{P}(\mathbb{P}^{2}(R)).$
@@
Two properly bounded left difference operators of degree three, $\mathcal{D},\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, are said to be related, denoted $\mathcal{D}\sim\widehat{\mathcal{D}}$, if
@@
In particular, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then, for any $\alpha\in\mathrm{Ker}\,\mathcal{D}$, we have $\widehat{\mathcal{D}}(\mathcal{D}_{+}\alpha)=\widehat{\mathcal{D}}_{+}(\mathcal{D}\alpha)=0.\qed$
@@
We say that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is regular if there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, and the homomorphism of right $R$-modules given by
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is regular. Then
@@
    The operator $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ exists and is unique up to the left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$. (Note that, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}\sim\alpha\widehat{\mathcal{D}}$ for any $\alpha\in(R^{\mathbb{Z}})^{*}$. Therefore, the relation $\mathcal{D}\sim\widehat{\mathcal{D}}$ is invariant under the left action of $(R^{\mathbb{Z}})^{*}$ on $\widehat{\mathcal{D}}$.)
@@
    Any operator in the same orbit of the left-right $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$ action on $R^{\mathbb{Z}}[T]_{3}^{pb}$ as $\mathcal{D}$ is also regular. Moreover, for two regular operators in the same orbit, their related operators are also in the same orbit.
@@
1. Existence of $\widehat{\mathcal{D}}$ with the property $\mathcal{D}\sim\widehat{\mathcal{D}}$ is a part of the definition of a regular operator. That $\widehat{\mathcal{D}}$ also satisfies $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For any other $\widehat{\mathcal{D}}^{\prime}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, we have $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\supset\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. Furthermore, we know that the mapping $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\to R^{3}$, given by taking three consecutive entries of a bi-infinite sequence, is an isomorphism. The same holds for $\widehat{\mathcal{D}}$. Therefore, the embedding $\mathrm{Ker}\,\widehat{\mathcal{D}}\to\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}$ must be an isomorphism as well. So, by Lemma 5.11, we have $\widehat{\mathcal{D}}^{\prime}=\alpha\widehat{\mathcal{D}}$ for some $\alpha\in(R^{\mathbb{Z}})^{*}$.
@@
As a result, for a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, all its related operators constitute a single orbit of the $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$ action. Moreover, this orbit remains the same if we replace $\mathcal{D}$ by a left-right-equivalent operator. Denote by $R^{\mathbb{Z}}[T]_{3}^{reg}$ the set of regular operators. Then we have a map
$\psi\colon R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$
which takes an orbit of a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ to the orbit of its related operator $\widehat{\mathcal{D}}\sim\mathcal{D}$ (this orbit is independent on the choice of a related operator, by Proposition 6.4).
Recall also that, by Theorem 5.7, we have a bijection $\varphi\colon R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the correspondence between difference operators (up to left-right equivalence) and polygons (up to projective equivalence).
@@
    $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))=\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}.$ (6.3)
@@
    ${R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})}$${R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})}$${\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}}$${\mathfrak{P}(\mathbb{P}^{2}(R)).}$$\scriptstyle{\psi}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{\Psi}$ (6.4)
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ is regular. We will show that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$. This will establish the inclusion $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))\subset\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$, as well as commutativity of the diagram (6.4). Let $v_{i}:=\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Then, by construction of the map $\bar{\varphi}\colon R^{\mathbb{Z}}[T]_{3}^{pb}\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the polygon $(Rv_{i})$ represents the class of polygons $\bar{\varphi}(\mathcal{D})$ associated to $\mathcal{D}$. Furthermore, by Proposition 5.8, we have that $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. Therefore,
@@
where $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is any operator such that $\mathcal{D}\sim\widehat{\mathcal{D}}$. So, if we set $\hat{v}:=\mathcal{D}_{+}v$, by Proposition 5.6 we have that $\hat{v}_{i},\hat{v}_{i+1},\hat{v}_{i+2}$ is a basis in $(\mathrm{Ker}\,\widehat{\mathcal{D}})^{*}$ for all $i$. In other words, the points $R\hat{v}_{i}$ form a polygon. At the same time, if $\mathcal{D}=a+bT+cT^{2}+dT^{3}$, then
@@
So, $\hat{v}_{i}\in(Rv_{i}\oplus Rv_{i+1})\cap(Rv_{i+2}\oplus Rv_{i+3})$, meaning that the point $R\hat{v}_{i}$ lies on the lines $Rv_{i}\oplus Rv_{i+1}$, $Rv_{i+2}\oplus Rv_{i+3}$ (those are indeed lines by Lemma 4.6) and hence coincides with their intersection point (which is unique also by Lemma 4.6). In other words, the points $R\hat{v}_{i}$ are the vertices of the image of the polygon $(Rv_{i})$ under the inverse pentagram map, which proves that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$.
Now, we show that $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))\supset\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$. Take a polygon $(Rv_{i})$, $v_{i}\in M$ (where $M$ is a rank three free left $R$-module), such that its image under the inverse pentagram map is a polygon. We need to show that the corresponding difference operator $\mathcal{D}$ satisfying $\mathcal{D}v=0$ is regular. Arguing as above, we see that the image of the polygon $(Rv_{i})$ under the inverse pentagram map is given by the points $R\hat{v}_{i}$, where $\hat{v}:=\mathcal{D}_{+}v$. Since this image is a polygon, there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\widehat{\mathcal{D}}\hat{v}=0$. Let’s show that $\mathcal{D}\sim\widehat{\mathcal{D}}$ and, furthermore, $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For the latter, note that
@@
$\tilde{\mathcal{D}}:=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}-\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}.$

### v3

@@ 5.2 The correspondence between difference operators and polygons
Let $R$ be stably finite. Suppose $v_{i}\in M$ are such that the sequence $P_{i}=Rv_{i}$ is a polygon. Then there exists a properly bounded degree $d$ difference operator $\mathcal{D}{{}_{v}}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ such that $\mathcal{D}{{}_{v}}v=0$.
@@
The following says that the correspondence $v\mapsto\mathcal{D}_{v}$ from Proposition 5.13 provides a right inverse for the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Thus, this map is surjective, and the proof of Theorem 5.7 is complete.
@@
Assume that $v_{i}\in M$, where $M$ is a rank $d$ free left $R$-module for stably finite ring $R$, is such that the sequence $Rv_{i}$ is a polygon. Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ be a properly bounded degree $d$ difference operator such that $\mathcal{D}v=0$. Then there is a projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ mapping the polygon associated with $\mathcal{D}$ to the polygon $(Rv_{i})$.
@@
Since $\mathcal{D}v=0$, and the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$, by Proposition 5.6 we have $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. This means that the image of the mapping $\psi\colon M^{*}\to R^{\mathbb{Z}}$ given by $\xi\mapsto(\xi(v_{i}))$ is precisely $\mathrm{Ker}\,\mathcal{D}$. Furthermore, since $M^{*}$ and $\mathrm{Ker}\,\mathcal{D}$ are free modules of the same rank, and $R$ is stably finite, it follows that $\psi\colon M^{*}\to\mathrm{Ker}\,\mathcal{D}$ is a right $R$-module isomorphism. Consider the dual $\psi^{*}\colon(\mathrm{Ker}\,\mathcal{D})^{*}\to M$. It is a left $R$-module isomorphism, and, for any $\xi\in M^{*}$, we have
@@ 6 Integrability of pentagram maps over rings
Throughout this section, the ground ring $R$ is assumed to be stably finite.
@@ 6.1 Pentagram maps in terms of difference operators
By Theorem 5.7, polygons in the (left) projective plane $\mathbb{P}^{2}(R)$, considered up to projective transformation, are in one-to-one correspondence with elements of the set $R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$ of properly bounded left difference operators of degree three, up to left-right action of $(R^{\mathbb{Z}})^{\times}$. Accordingly, one can describe the pentagram map as a dynamical system on difference operators. In this section we give an intrinsic description of that dynamical system. For technical reasons, we consider the inverse pentagram map, rather than the pentagram map itself. It turns out that that this inverse, phrased in terms of difference operators, is described by a simple linear equation. In the case when the ground ring is a field, this result was established in [9, Theorem 1.1].
In order to have well-defined dynamics on difference operators, we only consider polygons whose image under the inverse pentagram map is a polygon (so that the image corresponds to a difference operator). Denote the inverse pentagram map by $\Psi$, and let $\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}$ be the set of projective equivalence classes of polygons in $\mathbb{P}^{2}(R)$ whose image under the inverse pentagram map is a polygon. That way, we have a map
$\Psi\colon\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}\to\mathfrak{P}(\mathbb{P}^{2}(R)).$
@@
Two properly bounded left difference operators of degree three, $\mathcal{D},\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$, are said to be related, denoted $\mathcal{D}\sim\widehat{\mathcal{D}}$, if
@@
In particular, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then, for any $\alpha\in\mathrm{Ker}\,\mathcal{D}$, we have $\widehat{\mathcal{D}}(\mathcal{D}_{+}\alpha)=\widehat{\mathcal{D}}_{+}(\mathcal{D}\alpha)=0.$
@@
We say that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ is regular if there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, and the homomorphism of right $R$-modules given by
@@
When $R$ is the field of real or complex numbers, the set of $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ satisfying the first condition of this definition (existence of $\widehat{\mathcal{D}}\sim\mathcal{D}$) is dense by [9, Lemma 4.10]. Moreover, surjectivity of $\mathcal{D}_{+}\mid_{\mathrm{Ker}\,\mathcal{D}}$ is, in this case, equivalent to the requirement $\mathrm{Ker}\,\mathcal{D}_{+}\cap\mathrm{Ker}\,D_{-}=0$, which also holds on a dense subset. So, for real or complex numbers, regular operators are dense, hence the name.
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ is regular. Then
@@
    The operator $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ exists and is unique up to the left action of $(R^{\mathbb{Z}})^{\times}$ on $R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$. (Note that, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}\sim\alpha\widehat{\mathcal{D}}$ for any $\alpha\in(R^{\mathbb{Z}})^{\times}$. Therefore, the relation $\mathcal{D}\sim\widehat{\mathcal{D}}$ is invariant under the left action of $(R^{\mathbb{Z}})^{\times}$ on $\widehat{\mathcal{D}}$.)
@@
    Any operator in the same orbit of the left-right $(R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times}$ action on $R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ as $\mathcal{D}$ is also regular. Moreover, for two regular operators in the same orbit, their related operators are also in the same orbit.
@@
1. Existence of $\widehat{\mathcal{D}}$ with the property $\mathcal{D}\sim\widehat{\mathcal{D}}$ is a part of the definition of a regular operator. That $\widehat{\mathcal{D}}$ also satisfies $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For any other $\widehat{\mathcal{D}}^{\prime}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, we have $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\supset\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. Furthermore, we know that the mapping $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\to R^{3}$, given by taking three consecutive entries of a bi-infinite sequence, is an isomorphism. The same holds for $\widehat{\mathcal{D}}$. Therefore, the embedding $\mathrm{Ker}\,\widehat{\mathcal{D}}\to\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}$ must be an isomorphism as well. By Lemma 5.11, we have $\widehat{\mathcal{D}}^{\prime}=\alpha\widehat{\mathcal{D}}$ for some $\alpha\in(R^{\mathbb{Z}})^{\times}$.
@@
As a result, for a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$, all its related operators constitute a single orbit of the $(R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times}$ action. Moreover, this orbit remains the same if we replace $\mathcal{D}$ by a left-right-equivalent operator. Denote by $R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}$ the set of regular operators. Then we have a map
$\psi\colon R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})\to R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$
which takes an orbit of a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}$ to the orbit of its related operator $\widehat{\mathcal{D}}\sim\mathcal{D}$. This orbit is independent of the choice of a related operator, by Proposition 6.4.
Recall also that, by Theorem 5.7, we have a bijection $\varphi\colon R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the correspondence between difference operators (up to left-right equivalence) and polygons (up to projective equivalence).
@@
    $\varphi(R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times}))=\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}.$ (6.3)
@@
    $R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$$R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$$\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}$$\mathfrak{P}(\mathbb{P}^{2}(R)).$$\scriptstyle{\psi}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{\Psi}$ (6.4)
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}$. We will show that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$. This will establish the inclusion $\varphi(R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times}))\subset\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}$, as well as commutativity of the diagram (6.4). Let $v_{i}:=\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Then, by construction of the map $\bar{\varphi}\colon R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the polygon $(Rv_{i})$ represents the class of polygons $\bar{\varphi}(\mathcal{D})$ associated to $\mathcal{D}$. Furthermore, by Proposition 5.8, we have that $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. Therefore,
@@
[…]
@@
Thus, $\hat{v}_{i}\in(Rv_{i}\oplus Rv_{i+1})\cap(Rv_{i+2}\oplus Rv_{i+3})$, meaning that the point $R\hat{v}_{i}$ lies on the lines $Rv_{i}\oplus Rv_{i+1}$, $Rv_{i+2}\oplus Rv_{i+3}$ (those are indeed lines by Lemma 4.6) and hence coincides with their intersection point (which is unique also by Lemma 4.6). In other words, the points $R\hat{v}_{i}$ are the vertices of the image of the polygon $(Rv_{i})$ under the inverse pentagram map, which proves that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$.
Now, we show that $\varphi(R^{\mathbb{Z}}[T]_{3}^{\mathrm{reg}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times}))\supset\mathfrak{P}(\mathbb{P}^{2}(R))^{\mathrm{reg}}$. Take a polygon $(Rv_{i})$, $v_{i}\in M$ (where $M$ is a rank three free left $R$-module), such that its image under the inverse pentagram map is a polygon. We need to show that the corresponding difference operator $\mathcal{D}$ satisfying $\mathcal{D}v=0$ is regular. Arguing as above, we see that the image of the polygon $(Rv_{i})$ under the inverse pentagram map is given by the points $R\hat{v}_{i}$, where $\hat{v}:=\mathcal{D}_{+}v$. Since this image is a polygon, there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{\mathrm{pb}}$ such that $\widehat{\mathcal{D}}\hat{v}=0$. Let us show that $\mathcal{D}\sim\widehat{\mathcal{D}}$ and, furthermore, $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For the latter, note that
@@
$\tilde{\mathcal{D}}:=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}-\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}.$ (6.5)
@@
In [9], similar ideas are used to work with the pentagram map itself instead of its inverse. The former is also described by (6.2), but with $\mathcal{D}_{\pm}$ of the form
$\mathcal{D}_{+}:=a+cT^{2},\quad\mathcal{D}_{-}:=bT+dT^{3}.$ (6.6)
In that setting, Theorem 6.5 does not seem to be true as stated. The argument breaks down where we say that the operator $\tilde{\mathcal{D}}:=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}-\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}$ is of degree two; that is no longer the case for operators of the form (6.6).